The total surface area of a cylinder of height 9.5 cm is estimated using the function f(r) = 6r2 + 60r, where r is the radius of the cylinder. If the radius of the cylinder is 5 cm, what is its total surface area?

answers:

a) 1200 cm(squared)
b) 450 cm(squared)
c) 187.5 cm(squared)
d) 60 cm(squared)

Answer by stanbon(75887)

The total surface area of a cylinder of height 9.5 cm is estimated using the function f(r) = 6r2 + 60r, where r is the radius of the cylinder. If the radius of the cylinder is 5 cm, what is its total surface area?
--------
f(5) = 6*5^2+50*5
---
f(5) = 150+250 = 400 cm^2
----------------------------
That's f(5). It doesn't show any relation
to the 9.5 cm height which certainly is
a factor in determining the total surface
area.
Cheers,
Stan H.
========
answers:
a) 1200 cm(squared)
b) 450 cm(squared)
c) 187.5 cm(squared)
d) 60 cm(squared)